OpenStudy (anonymous):

The population of a local species of beetle can be found using an infinite geometric series where a1 = 960 and the common ratio is one forth. Write the sum in sigma notation, and calculate the sum (if possible) that will be the upper limit of this population. the summation of 960 times one fourth to the i minus 1 power, from i equals 1 to infinity. ; the sum is 1,280 the summation of 960 times one fourth to the i minus 1 power, from i equals 1 to infinity. ; the sum is divergent the summation of 960 times one fourth to the i power, from i equals 1 to infinity. ; the sum is 1,280

OpenStudy (anonymous):

the summation of 960 times one fourth to the i power, from i equals 1 to infinity. ; the series is divergent

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

I need help on this @ganeshie8

OpenStudy (anonymous):

Please @ganeshie8

OpenStudy (anonymous):

At least can I tell what I would respond @ganeshie8

ganeshie8 (ganeshie8):

\[\large \sum \limits_{i=1}^{\infty} 960 \left(\frac{1}{4}\right)^{i-1}\]

ganeshie8 (ganeshie8):

familiar with infinite geometric series formula ?

OpenStudy (anonymous):

Yes @ganeshie8

ganeshie8 (ganeshie8):

use it and find the infinite sum

ganeshie8 (ganeshie8):

\[\large S_{\infty} = \dfrac{a_1}{1-r}\]

ganeshie8 (ganeshie8):

\(a_1\) = 960 \(r\) = 1/4 plugin and simplify

OpenStudy (anonymous):

1280

ganeshie8 (ganeshie8):

yes!

OpenStudy (anonymous):

Ok so the answer would be A or C

ganeshie8 (ganeshie8):

A